OpenStudy (anonymous):

A charged particle is projected in the air such that its horizontal (x) and vertical (y) shifts are given by x(t) = 10t − t2 and y(t) = 6t, where x and y are in meters and time t is in seconds. What is the ratio of y to x at t = 4 seconds?

OpenStudy (whpalmer4):

\[y(t):x(t) = \frac{6t}{10t-t^2}\]Evaluate that at \(t = 4\)

OpenStudy (anonymous):

A. 0.25 B. 2 C. 0.75 D. 0.85 E. 1

OpenStudy (whpalmer4):

Yeah, it's one of those. Do the arithmetic.

OpenStudy (anonymous):

dont have a calculator

OpenStudy (whpalmer4):

Don't need one. What's 6*4?

OpenStudy (whpalmer4):

What's 10*4 What's 4*4 What's \[\frac{6*4}{(10*4)-(4*4)}\]

OpenStudy (anonymous):

24

OpenStudy (whpalmer4):

No. \[\frac{6*4}{(10*4)-(4*4)}= \frac{24}{40-16} = \frac{24}{24} =? \]